# On the magnetic perturbation theory for Chern insulators

### v1

@@ 2.1. Some heuristics behind the gauge covariant magnetic perturbation theory
we have that $U_{\theta}(-\mathrm{i}\nabla_{\bf x}-A^{\prime})^{2}U_{\theta}^{-1}=(-\mathrm{i}\nabla_{\bf x}-A)^{2}$ It can be shown that the operator $U_{\theta}$ actually provides a unitary equivalence between $H^{2}_{A^{\prime}}({\mathbb{R}}^{2})$ and $H^{2}_{A}({\mathbb{R}}^{2})$, thus showing that $(-\mathrm{i}\nabla_{\bf x}-A^{\prime})^{2}$ and $(-\mathrm{i}\nabla_{\bf x}-A)^{2}$ are unitarily equivalent. $U_{\theta}$ is called gauge transform and the function $\theta$ is usually called the generator of the gauge transformation. Recall that the key formula (2.9) is nothing but the gauge covariance of the magnetic momentum $(-\mathrm{i}\nabla-A)$ under the gauge transformation generated by $\theta({\bf x})=\varphi({\bf x},{\bf x}^{\prime})$.
@@ 3. Approximate perturbed projections
Let us show that both $P_{b}$ and $P_{b_{0}}$ are integral operators with a jointly continuous integral kernel. We need a general technical lemma from [43, Lemma B.7.8]:
@@
Let $C$ be a bounded operator on $L^{2}({\mathbb{R}}^{2})$ with an integral kernel $C:{\mathbb{R}}^{2}\times{\mathbb{R}}^{2}\to{\mathbb{C}}$ with the property that the map ${\mathbb{R}}^{2}\ni{\bf y}\mapsto C(\cdot;{\bf y})\in L^{2}({\mathbb{R}}^{2})$ is continuous.
@@
Let us define the $L^{2}$ functions $C_{\bf y}:=C(\cdot;{\bf y})$ and $A^{*}_{\bf x}:=\overline{A({\bf x};\cdot)}$. Then we can put
$D({\bf x};{\bf y}):=\langle A^{*}_{\bf x},BC_{\bf y}\rangle\,.$
@@
$|P_{b}({\bf x};{\bf y})|\leq C{\mathrm{e}}^{-\alpha\|{\bf x}-{\bf y}\|}\,,\quad\forall\,{\bf x},{\bf y}\in{\mathbb{R}}^{2}.$ (3.1)
@@
$\widetilde{P}_{b}({\bf x};{\bf y}):=e^{\mathrm{i}\epsilon\varphi({\bf x},{\bf y})}P_{b_{0}}({\bf x};{\bf y})\,.$ (3.2)
@@
$P_{b}=\widetilde{P}_{b}+\frac{\mathrm{i}}{2\pi}\sum_{n=1}^{\infty}\oint_{\mathcal{C}}(-1)^{n}\,S_{\epsilon}(z)(T_{\epsilon}(z))^{n}\,\mathrm{d}z\,,$ (3.3)
@@
$|P_{b}({\bf x};{\bf y})-\widetilde{P}_{b}({\bf x};{\bf y})|\leq|\epsilon|\,C{\mathrm{e}}^{-\alpha\|{\bf x}-{\bf y}\|}\,,\quad\|P_{b}-\widetilde{P}_{b}\|\leq|\epsilon|\,C\,.$ (3.4)
@@
$|\Delta^{(b)}({\bf x};{\bf x}^{\prime})|\leq|\epsilon|\,Ce^{-\alpha\|{\bf x}-{\bf x}^{\prime}\|}\,.$ (3.6)
Via a Schur test [26, Theorem 5.2] , we conclude that $\|\Delta^{(b)}\|\leq|\epsilon|\,C$. The operator $\Delta^{(b)}$ is a function of the selfadjoint operator $\widetilde{P}_{b}$ with a norm of order $|\epsilon|$. By functional calculus, the spectrum of $\widetilde{P}_{b}$ needs to be concentrated near 00 or $1$, which is compatible with the fact that $\widetilde{P}_{b}$ is an almost projection.
@@
$\mathcal{P}^{(b)}:=\widetilde{P}_{b}+\big{(}\widetilde{P}_{b}-\tfrac{1}{2}{\bf 1}\big{)}\big{\{}({\bf 1}+4\Delta^{(b)})^{-1/2}-{\bf 1}\big{\}}\,$ (3.7)
@@
$\big{|}\mathcal{P}^{(b)}({\bf x};{\bf y})-\widetilde{P}_{b}({\bf x};{\bf y})\big{|}\leq C\,|\epsilon|\,e^{-\alpha^{\prime}\|{\bf x}-{\bf y}\|}.$ (3.8)
Moreover, let $\chi_{L}$ be the characteristic function of the set $\Lambda_{L}:=[-L,L]^{2}$, $\chi_{L}P_{b}$ and $\chi_{L}\mathcal{P}^{(b)}$ are trace class and
$\lim_{L\to\infty}\left|\frac{\operatorname{Tr}(\chi_{L}P_{b})}{4L^{2}}-\frac{\operatorname{Tr}(\chi_{L}\mathcal{P}^{(b)})}{4L^{2}}\right|=0\,.$ (3.9)
@@
$(\widetilde{P}_{b}-z\mathbf{1})^{-1}$ $=\frac{\widetilde{P}_{b}+(z-1)\mathbf{1}}{z(1-z)}\left({\bf 1}+\frac{\Delta^{(b)}}{z(1-z)}\right)^{-1}$ (3.10)
@@
$\left(\Delta^{(b)}-z(z-1){\bf 1}\right)$ $=-\left(\frac{{\bf 1}+({\bf 1}+4\Delta^{(b)})^{1/2}}{2}-z{\bf 1}\right)\left(\frac{{\bf 1}-({\bf 1}+4\Delta^{(b)})^{1/2}}{2}-z{\bf 1}\right)$ (3.11)
@@
For $\epsilon$ close to zero, the norm of $\Delta^{(b)}$ is also close to zero, which implies, by functional calculus, that the spectrum of $C^{(b)}$ is concentrated near $1$ and the spectrum of $D^{(b)}$ is concentrated near 00. By inserting (3.11) in (3.10) and using the second resolvent identity, we get that $\mathcal{P}^{(b)}$ can be expressed as
@@
$\frac{\mathrm{i}}{2\pi}\oint_{\mathcal{C}_{1}}\,\left(C^{(b)}-z{\bf 1}\right)^{-1}\mathrm{d}z={\bf 1}$
@@
Now we want to prove (3.8). In order to analyze the integral kernel of $\mathcal{P}^{(b)}$, we first show a technical lemma, which is a generalization of [16, Lemma 8.5].
@@
$\left|D_{\epsilon}({\bf x};{\bf y})\right|\leq|\epsilon|\,C_{0}e^{-\alpha_{0}\|{\bf x}-{\bf y}\|}\,\,\,\,\,\forall\,{\bf x}\,,{\bf y}\in{\mathbb{R}}^{2}\,,\,\,\text{ and }\,\,\epsilon_{0}\,C_{0}\int_{{\mathbb{R}}^{2}}e^{-\alpha_{0}\|{\bf x}\|/2}\,\mathrm{d}{\bf x}<r.$ (3.12)
@@
Due to (3.12) and the Schur test [26, Theorem 5.2] we have $\|D_{\epsilon}\|<r$. The operator $D_{\epsilon}^{n}$ with $n\geq 2$ has an integral kernel given by:
@@
${\mathrm{e}}^{\alpha_{0}\|{\bf x}-{\bf x}^{\prime}\|/2}\left|D_{\epsilon}^{n}({\bf x};{\bf x}^{\prime})\right|\leq|\epsilon|\,C_{0}\,\,r^{n-1}.$ (3.13)
@@
$\left|\left(P_{b}-\mathcal{P}^{(b)}\right)({\bf x};{\bf x}^{\prime})\right|=\left|\left(P_{b}-\widetilde{P}_{b}+\widetilde{P}_{b}-\mathcal{P}^{(b)}\right)({\bf x};{\bf x}^{\prime})\right|\leq|\epsilon|\,C\,e^{-\alpha\|{\bf x}-{\bf x}^{\prime}\|}\,.$ (3.14)
Due to (3.14) we have that $\|P_{b}-\mathcal{P}^{(b)}\|<1$ when $\epsilon$ is sufficiently small. Since the norm difference of the two projections, $P_{b}$ and $\mathcal{P}^{(b)}$, is strictly less than one, we can construct the Kato-Nagy operator $\mathsf{U}_{b}$ [27, Sec. 1.6.8](see (3.15) for the explicit expression), which is unitary and intertwines the two projections, namely $\mathsf{U}_{b}P_{b}=\mathcal{P}^{(b)}\mathsf{U}_{b}$. The next Lemma is adapted from [16, Lemma 8.5].
@@
$\mathsf{U}_{b}$ $=\left[{\bf 1}-\left({\mathcal{P}}^{(b)}-P_{b}\right)^{2}\right]^{-1/2}\left({\mathcal{P}}^{(b)}P_{b}+\big{(}{\bf 1}-{\mathcal{P}}^{(b)}\big{)}\left({\bf 1}-P_{b}\right)\right).$ (3.15)
@@
$\left|\left(\mathsf{U}_{b}-{\bf 1}\right)({\bf x};{\bf x}^{\prime})\right|\leq|\epsilon|\,C\,{\mathrm{e}}^{-\alpha\|{\bf x}-{\bf x}^{\prime}\|}\,,\qquad\forall\,{\bf x}\,,{\bf y}\in{\mathbb{R}}^{2}.$ (3.16)
@@
The function $F(x)=(1-x^{2})^{-1/2}-1$ is real analytic on $(-1,1)$ and due to (3.14) and Lemma 3.5 we conclude that if $\epsilon_{0}$ is small enough, then $F\big{(}{\mathcal{P}}^{(b)}-P_{b}\big{)}$ has an integral kernel of order $\epsilon$ with an off-diagonal exponential decay. By using that $P_{b}^{2}+({\bf 1}-P_{b})^{2}={\bf 1}$, we observe that $\mathsf{U}_{b}-{\bf 1}$ can be expressed as a sum of products involving at least one factor like $F\big{(}{\mathcal{P}}^{(b)}-P_{b}\big{)}$ or ${\mathcal{P}}^{(b)}-P_{b}$, and we are done with proving (3.16). ∎
@@
$\operatorname{Tr}(\chi_{L}P_{b})-\operatorname{Tr}(\chi_{L}\mathcal{P}^{(b)})=\mathcal{O}(L)$
and (3.9) immediately follows. ∎
4. Smoothness in the strong topology and stability of the discrete spectrum
@@
$\big{|}(\widetilde{P}_{b}-P_{b_{0}})({\bf x};{\bf y})\big{|}\leq\big{|}e^{\mathrm{i}\epsilon\varphi({\bf x},{\bf y})}-1\big{|}\,|P_{b_{0}}({\bf x};{\bf y})|\leq C\,|\epsilon|\,\|{\bf x}\|\,\|{\bf y}\|\,e^{-\alpha\|{\bf x}-{\bf y}\|}.$ (4.1)
@@
$\sup_{z\in\mathcal{C}}\|W_{\epsilon}(H_{b_{0}}-z{\bf 1})^{-1}e^{-\alpha\|\cdot\|}\|\leq C\,|\epsilon|$ (4.2)
@@
$\sup_{z\in\mathcal{C}}\|\big{(}(H_{b}-z{\bf 1})^{-1}-(H_{b_{0}}-z{\bf 1})^{-1}\big{)}e^{-\alpha\|\cdot\|}\|\leq C\,|\epsilon|\,.$ (4.3)
@@
$\sup_{{\rm dist}(\mu,\mathcal{C})\geq\delta>0}\|\big{(}R_{b}(\mu)-R_{b_{0}}(\mu)\big{)}\,e^{-\alpha\|\cdot\|}\|\leq C\,|\epsilon|.$ (4.4)
@@ 4.1. Perturbation of discrete eigenvalues
Let us now consider the situation where the spectral island $\sigma_{0}$, enclosed by $\mathcal{C}$, consists of an isolated discrete eigenvalue $e_{0}$ of $H_{b_{0}}$, with multiplicity $M\geq 1$.
From a result by Kato [27, VIII Theorem 2.6] (see also [1, Section §$\S$ 6]) we have that $\sigma_{b_{0}}$ will evolve into a spectral island $\sigma_{b}$ that contains $M$ discrete eigenvalues $e^{(i)}_{b}$ (including multiplicities). Let us show how this fact can be proved within magnetic perturbation theory.
@@
$P_{b_{0}}({\bf x};{\bf y})=\sum_{i=1}^{M}{\psi_{i}({\bf x})}\overline{\psi_{i}({\bf y})}\,,\quad|P_{b_{0}}({\bf x};{\bf y})|\leq C\,e^{-\alpha\|{\bf x}\|}e^{-\alpha\|{\bf y}\|},$
where $\psi_{i}$ forms a basis of the eigenspace associated to $e_{0}$. Then, we have the estimate
$\left|\left(\widetilde{P}_{b}-P_{b_{0}}\right)({\bf x};{\bf y})\right|$ $\leq\left|e^{\mathrm{i}\epsilon\varphi({\bf x},{\bf y})}-1\right|\left|P_{b_{0}}({\bf x};{\bf y})\right|$ (4.6)
@@
where we have used that $|\varphi({\bf x};{\bf y})|\leq C\,\|{\bf x}\|\|{\bf y}\|$. Then a Schur test [26, Theorem 5.2] finishes the proof. ∎
@@
$\mathfrak{W}_{\epsilon}({\bf x};{\bf x}^{\prime})=2e^{\mathrm{i}\epsilon\varphi({\bf x},{\bf x}^{\prime})}\frac{\mathrm{i}}{2\pi}\oint_{\mathcal{C}}\mathrm{d}z\int_{{\mathbb{R}}^{2}}(H_{b_{0}}-z{\bf 1})^{-1}({\bf x};{\bf y})$
@@
$\|\big{(}\widetilde{H}_{b}-\mu({\bf 1}-P_{b_{0}})\big{)}^{-1}-({\bf 1}-P_{b_{0}})\,R_{b}(\mu)({\bf 1}-P_{b_{0}})\,\|\leq C\,|\epsilon|.$ (4.9)
@@
$\Big{(}\widetilde{H}_{b}-\mu({\bf 1}-P_{b_{0}})\Big{)}\,({\bf 1}-P_{b_{0}})\,R_{b}(\mu)\,({\bf 1}-P_{b_{0}})$ (4.10)
$\qquad=({\bf 1}-P_{b_{0}})-[W_{\epsilon},P_{b_{0}}]R_{b}(\mu)({\bf 1}-P_{b_{0}})+\mathcal{O}(\epsilon)$
@@
$\mu=e_{0}+\langle\psi,W_{\epsilon}\psi\rangle-\langle W_{\epsilon}\psi,\big{(}\widetilde{H}_{b}-\mu({\bf 1}-P_{b_{0}})\big{)}^{-1}W_{\epsilon}\psi\rangle.$ (4.11)
This is nothing but a fixed-point equation whose iterations starting from $\mu=e_{0}$ gives us an asymptotic expansion in $\epsilon$ of $e_{\epsilon}$. We a-priori know that $e_{\epsilon}=e_{0}+\mathcal{O}(\epsilon)$ and $W_{\epsilon}\psi=\mathcal{O}(\epsilon)$, hence using (4.9) and (4.4) we obtain

### v2

@@ 2.1. Some heuristics behind the gauge covariant magnetic perturbation theory
we have that $U_{\theta}(-\mathrm{i}\nabla_{\bf x}-A^{\prime})^{2}U_{\theta}^{-1}=(-\mathrm{i}\nabla_{\bf x}-A)^{2}$ . It can be shown that the operator $U_{\theta}$ actually provides a unitary equivalence between $H^{2}_{A^{\prime}}({\mathbb{R}}^{2})$ and $H^{2}_{A}({\mathbb{R}}^{2})$, thus showing that $(-\mathrm{i}\nabla_{\bf x}-A^{\prime})^{2}$ and $(-\mathrm{i}\nabla_{\bf x}-A)^{2}$ are unitarily equivalent. $U_{\theta}$ is called gauge transform and the function $\theta$ is usually called the generator of the gauge transform. Recall that the key formula (2.7) is nothing but the gauge covariance of the magnetic momentum $(-\mathrm{i}\nabla-A)$ under the gauge transform generated by $\theta({\bf x})=\varphi({\bf x},{\bf x}^{\prime})$.
@@ 3. Approximate perturbed projections
Let us show that both $P_{b}$ and $P_{b_{0}}$ are integral operators with a jointly continuous integral kernel. We need a general technical lemma from [44, Lemma B.7.8]:
@@
Let $C$ be a bounded operator on $L^{2}({\mathbb{R}}^{2})$ with an integral kernel $C:{\mathbb{R}}^{2}\times{\mathbb{R}}^{2}\to{\mathbb{C}}$ with the property that the map ${\mathbb{R}}^{2}\ni{\bf y}\mapsto C(\cdot;{\bf y})\in L^{2}({\mathbb{R}}^{2})$ is continuous in the $L^{2}$ norm.
@@
Let us define the $L^{2}$ functions $C_{\bf y}:=C(\cdot;{\bf y})$ and $A^{*}_{\bf x}:=\overline{A({\bf x};\cdot)}$. Then it is enough to set $D({\bf x};{\bf y}):=\langle A^{*}_{\bf x},BC_{\bf y}\rangle\,.$
@@
(3.1) $|P_{b}({\bf x};{\bf y})|\leq Ce^{-\alpha\|{\bf x}-{\bf y}\|}\,,\quad\forall\,{\bf x},{\bf y}\in{\mathbb{R}}^{2}.$
@@
(3.2) $\widetilde{P}_{b}({\bf x};{\bf y}):=e^{\mathrm{i}\epsilon\varphi({\bf x},{\bf y})}P_{b_{0}}({\bf x};{\bf y})\,.$
@@
(3.3) $P_{b}=\widetilde{P}_{b}+\frac{\mathrm{i}}{2\pi}\sum_{n=1}^{\infty}\oint_{\mathcal{C}}(-1)^{n}\,S_{\epsilon}(z)(T_{\epsilon}(z))^{n}\,\mathrm{d}z\,$,
@@
(3.4) $|P_{b}({\bf x};{\bf y})-\widetilde{P}_{b}({\bf x};{\bf y})|\leq|\epsilon|\,Ce^{-\alpha\|{\bf x}-{\bf y}\|}\,,\quad\|P_{b}-\widetilde{P}_{b}\|\leq|\epsilon|\,C\,.$
@@
$|\Delta^{(b)}({\bf x};{\bf x}^{\prime})|\leq|\epsilon|\,Ce^{-\alpha\|{\bf x}-{\bf x}^{\prime}\|}\,.$
Via a Schur test [27, Theorem 5.2], we conclude that $\|\Delta^{(b)}\|\leq|\epsilon|\,C$. The operator $\Delta^{(b)}$ is a function of the selfadjoint operator $\widetilde{P}_{b}$ with a norm of order $|\epsilon|$. By functional calculus, the spectrum of $\widetilde{P}_{b}$ needs to be concentrated near 0 or $1$, which is compatible with the fact that $\widetilde{P}_{b}$ is an almost projection.
@@
(3.5) $\mathcal{P}^{(b)}:=\widetilde{P}_{b}+\big(\widetilde{P}_{b}-\tfrac{1}{2}{\bf 1}\big)\big\{({\bf 1}+4\Delta^{(b)})^{-1/2}-{\bf 1}\big\}\,$
@@
(3.6) $\big|\mathcal{P}^{(b)}({\bf x};{\bf y})-\widetilde{P}_{b}({\bf x};{\bf y})\big|\leq C\,|\epsilon|\,e^{-\alpha^{\prime}\|{\bf x}-{\bf y}\|}.$
Moreover, let $\chi_{L}$ be the characteristic function of the set $\Lambda_{L}:=[-L,L]^{2}$, then $\chi_{L}P_{b}$ and $\chi_{L}\mathcal{P}^{(b)}$ are trace class and
(3.7) $\lim_{L\to\infty}\left|\frac{\operatorname{Tr}\left(\chi_{L}P_{b}\right)}{4L^{2}}-\frac{\operatorname{Tr}(\chi_{L}\mathcal{P}^{(b)})}{4L^{2}}\right|=0\,.$
@@
(3.8) $(\widetilde{P}_{b}-z\mathbf{1})^{-1}$ $=\frac{\widetilde{P}_{b}+(z-1)\mathbf{1}}{z(1-z)}\left({\bf 1}+\frac{\Delta^{(b)}}{z(1-z)}\right)^{-1}$
@@
(3.9) $\left(\Delta^{(b)}-z(z-1){\bf 1}\right)$ $=-\left(\frac{{\bf 1}+({\bf 1}+4\Delta^{(b)})^{1/2}}{2}-z{\bf 1}\right)\left(\frac{{\bf 1}-({\bf 1}+4\Delta^{(b)})^{1/2}}{2}-z{\bf 1}\right)$
@@
For $\epsilon$ close to zero, the norm of $\Delta^{(b)}$ is also close to zero, which implies, by functional calculus, that the spectrum of $C^{(b)}$ is concentrated near $1$ and the spectrum of $D^{(b)}$ is concentrated near 0. By inserting (3.9) in (3.8) and using the second resolvent identity, we get that $\mathcal{P}^{(b)}$ can be expressed as
@@
$\frac{\mathrm{i}}{2\pi}\oint_{\mathcal{C}_{1}}\,\left(C^{(b)}-z{\bf 1}\right)^{-1}\mathrm{d}z={\bf 1}\,$,
@@
Now we want to prove (3.6). In order to analyze the integral kernel of $\mathcal{P}^{(b)}$, we first show a technical lemma, which is a generalization of [14, Lemma 8.5].
@@
(3.10) $\left|D_{\epsilon}({\bf x};{\bf y})\right|\leq|\epsilon|\,C_{0}e^{-\alpha_{0}\|{\bf x}-{\bf y}\|}\,\,\,\,\,\forall\,{\bf x}\,,{\bf y}\in{\mathbb{R}}^{2}\,,\,\,\text{ and }\,\,\epsilon_{0}\,C_{0}\int_{{\mathbb{R}}^{2}}e^{-\alpha_{0}\|{\bf x}\|/2}\,\mathrm{d}{\bf x}<r.$
@@
Due to (3.10) and the Schur test [27, Theorem 5.2] we have $\|D_{\epsilon}\|<r$. The operator $D_{\epsilon}^{n}$ with $n\geq 2$ has an integral kernel given by:
@@
$e^{\alpha_{0}\|{\bf x}-{\bf x}^{\prime}\|/2}\left|D_{\epsilon}^{n}({\bf x};{\bf x}^{\prime})\right|\leq|\epsilon|\,C_{0}\,\,r^{n-1}.$
@@
(3.11) $\left|\left(P_{b}-\mathcal{P}^{(b)}\right)({\bf x};{\bf x}^{\prime})\right|=\left|\left(P_{b}-\widetilde{P}_{b}+\widetilde{P}_{b}-\mathcal{P}^{(b)}\right)({\bf x};{\bf x}^{\prime})\right|\leq|\epsilon|\,C\,e^{-\alpha\|{\bf x}-{\bf x}^{\prime}\|}\,.$
Due to (3.11) we have that $\|P_{b}-\mathcal{P}^{(b)}\|<1$ when $\epsilon$ is sufficiently small. Since the norm difference of the two projections, $P_{b}$ and $\mathcal{P}^{(b)}$, is strictly less than one, we can construct the Kato-Nagy operator $\mathsf{U}_{b}$ [28, Sec. 1.6.8](see (3.12) for the explicit expression), which is unitary and intertwines the two projections, namely $\mathsf{U}_{b}P_{b}=\mathcal{P}^{(b)}\mathsf{U}_{b}$. The next Lemma is adapted from [14, Lemma 8.5].
@@
(3.12) $\mathsf{U}_{b}$ $=\left[{\bf 1}-\left({\mathcal{P}}^{(b)}-P_{b}\right)^{2}\right]^{-1/2}\left({\mathcal{P}}^{(b)}P_{b}+\big({\bf 1}-{\mathcal{P}}^{(b)}\big)\left({\bf 1}-P_{b}\right)\right).$
@@
(3.13) $\left|\left(\mathsf{U}_{b}-{\bf 1}\right)({\bf x};{\bf x}^{\prime})\right|\leq|\epsilon|\,C\,{\mathrm{e}}^{-\alpha\|{\bf x}-{\bf x}^{\prime}\|}\,,\qquad\forall\,{\bf x}\,,{\bf y}\in{\mathbb{R}}^{2}.$
@@
The function $F(x)=(1-x^{2})^{-1/2}-1$ is real analytic on (1,1)(-1,1) and due to (3.11) and Lemma 3.5 we conclude that if $\epsilon_{0}$ is small enough, then $F\big({\mathcal{P}}^{(b)}-P_{b}\big)$ has an integral kernel of order $\epsilon$ with an off-diagonal exponential decay. By using that $P_{b}^{2}+({\bf 1}-P_{b})^{2}={\bf 1}$, we observe that $\mathsf{U}_{b}-{\bf 1}$ can be expressed as a sum of products involving at least one factor like $F\big({\mathcal{P}}^{(b)}-P_{b}\big)$ or ${\mathcal{P}}^{(b)}-P_{b}$, and we are done with proving (3.13). ∎
@@
$\operatorname{Tr}(\chi_{L}P_{b})-\operatorname{Tr}(\chi_{L}\mathcal{P}^{(b)})=\mathcal{O}(L)\,$,
thus (3.7) immediately follows. ∎
4. Continuity in the strong topology and stability of the discrete spectrum
@@
$\big|(\widetilde{P}_{b}-P_{b_{0}})({\bf x};{\bf y})\big|\leq\big|e^{\mathrm{i}\epsilon\varphi({\bf x},{\bf y})}-1\big|\,|P_{b_{0}}({\bf x};{\bf y})|\leq C\,|\epsilon|\,\|{\bf x}\|\,\|{\bf y}\|\,e^{-\alpha\|{\bf x}-{\bf y}\|}.$
@@
(4.1) $\sup_{z\in\mathcal{C}}\left\|W_{\epsilon}(H_{b_{0}}-z{\bf 1})^{-1}e^{-\alpha\|\cdot\|}\right\|\leq C\,|\epsilon|$
@@
(4.2) $\sup_{z\in\mathcal{C}}\left\|\big((H_{b}-z{\bf 1})^{-1}-(H_{b_{0}}-z{\bf 1})^{-1}\big)e^{-\alpha\|\cdot\|}\right\|\leq C\,|\epsilon|\,.$
@@
(4.3) $\sup_{{\rm dist}(\mu,\mathcal{C})\geq\delta>0}\left\|\big(R_{b}(\mu)-R_{b_{0}}(\mu)\big)\,e^{-\alpha\|\cdot\|}\right\|\leq C\,|\epsilon|.$
@@ 4.1. Perturbation of discrete eigenvalues
Let us now consider the situation where the spectral island $\sigma_{0}\equiv\sigma_{b_{0}}$, enclosed by $\mathcal{C}$, consists of an isolated discrete eigenvalue $e_{0}$ of $H_{b_{0}}$, with multiplicity $M\geq 1$.
From a result by Kato [28, VIII Theorem 2.6] (see also [2, Section $\S$ 6]) we have that $\sigma_{b_{0}}$ will evolve into a spectral island $\sigma_{b}$ that contains $M$ discrete eigenvalues $e^{(i)}_{b}$ (including multiplicities). Let us show how this fact can be proved within magnetic perturbation theory.
@@
(4.4) $P_{b_{0}}({\bf x};{\bf y})=\sum_{i=1}^{M}{\psi_{i}({\bf x})}\overline{\psi_{i}({\bf y})}\,,\quad|P_{b_{0}}({\bf x};{\bf y})|\leq C\,e^{-\alpha\|{\bf x}\|}e^{-\alpha\|{\bf y}\|},$
where $\{\psi_{i}\}_{i=1}^{M}$ forms a basis of the eigenspace associated with $e_{0}$. Then, we have the estimate
$\left|\left(\widetilde{P}_{b}-P_{b_{0}}\right)({\bf x};{\bf y})\right|$ $\leq\left|e^{\mathrm{i}\epsilon\varphi({\bf x},{\bf y})}-1\right|\left|P_{b_{0}}({\bf x};{\bf y})\right|$
@@
where we have used that $|\varphi({\bf x},{\bf y})|\leq C\,\|{\bf x}\|\|{\bf y}\|$. Then a Schur test [27, Theorem 5.2] finishes the proof. ∎
@@
$\mathfrak{W}_{\epsilon}({\bf x};{\bf x}^{\prime})=2e^{\mathrm{i}\epsilon\varphi({\bf x},{\bf x}^{\prime})}\frac{\mathrm{i}}{2\pi}\oint_{\mathcal{C}}\int_{{\mathbb{R}}^{2}}(H_{b_{0}}-z{\bf 1})^{-1}({\bf x};{\bf y})$
@@
(4.7) $\left\|\big(\widetilde{H}_{b}-\mu({\bf 1}-P_{b_{0}})\big)^{-1}-({\bf 1}-P_{b_{0}})\,R_{b}(\mu)({\bf 1}-P_{b_{0}})\,\right\|\leq C\,|\epsilon|.$
@@
(4.8) $\Big(\widetilde{H}_{b}-\mu({\bf 1}-P_{b_{0}})\Big)\,({\bf 1}-P_{b_{0}})\,R_{b}(\mu)\,({\bf 1}-P_{b_{0}})$
$\qquad=({\bf 1}-P_{b_{0}})-({\bf 1}-P_{b_{0}})[W_{\epsilon},P_{b_{0}}]R_{b}(\mu)({\bf 1}-P_{b_{0}})+\mathcal{O}(\epsilon)$
@@
(4.9) $\mu=e_{0}+\langle\psi,W_{\epsilon}\psi\rangle-\langle W_{\epsilon}\psi,\big(\widetilde{H}_{b}-\mu({\bf 1}-P_{b_{0}})\big)^{-1}W_{\epsilon}\psi\rangle.$
This is nothing but a fixed-point equation whose iterations starting from $\mu=e_{0}$ give us an asymptotic expansion in $\epsilon$ of $e_{\epsilon}$. We a-priori know that $e_{\epsilon}=e_{0}+\mathcal{O}(\epsilon)$ and $W_{\epsilon}\psi=\mathcal{O}(\epsilon)$, hence using (4.7) and (4.3), together with the exponential localization of the eigenfunction (see (4.4)) and (1.5), we obtain